OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. -4 + yi = x + 3i

hartnn (hartnn):

just compare the real and imaginary terms! like if a+ib = m+in then a=m b=n

OpenStudy (anonymous):

you should compare imaginary and real form :)

OpenStudy (anonymous):

-4=x yi=3i

hartnn (hartnn):

yi=3i gives y=3 isn't it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is the answer i got i just wanted to make sure it was correct (:

hartnn (hartnn):

you're smart ;) you'll get it correct :)